OpenStudy (jewelotaku2001):

Which value for x makes the open sentence true? 7 + 6 • x = 4^2 + x^2 a) 5 b) 4 c) 3 d) 2

OpenStudy (danjs):

Change it to standard form first ax^2 + bx + c = 0 x^2 - 6x + 9 = 0 can you factor that now

OpenStudy (jewelotaku2001):

@AloneS

OpenStudy (skullpatrol):

Recall the zero-product property ab=0 if and only if a=0 or b=0.

OpenStudy (jewelotaku2001):

But there is no a or b.... there's x. I'm confused.

OpenStudy (skullpatrol):

The "a or b" stands for any variable :-)

OpenStudy (skullpatrol):

Do you see how DanJS got x^2 -6x +9 = 0?

OpenStudy (jewelotaku2001):

Nope, no idea.

OpenStudy (skullpatrol):

$$7 + 6 • x = 4^2 + x^2$$ $$7 + 6 • x -7-6•x= 4^2 + x^2-7-6•x$$

OpenStudy (skullpatrol):

Subtract 7 and 6•x from both sides.

OpenStudy (jewelotaku2001):

So then, on the left side of the equal sign, would it be x^2?

OpenStudy (skullpatrol):

Nope, it would be 0.

OpenStudy (jewelotaku2001):

Alright I think I can figure it out now. Thanks.

OpenStudy (skullpatrol):

Thanks for trying :-)